What is the probability that if 14 letters are typed, no letters are repeated?
Answer #1

There are Total 14 Letters are typed  and There are 26 letters in alphabets

So therefore Total possible letters = N_Total = 26^14

And We have to type without repetition

So in first place 26 possibilities may come because there are 26 letters

In second 25 letters because one letter already occupy the place And So On we have to type 14 letters

So there Possible letter words without repetition is = 26*25*24*23*22*21*20*19*18*17*16*15*14*13

so all the 14 letters are came without repetition.

Probability that if 14 letters are typed, no letters are repeated = (26^14) / (26*25*24*23*22*21*20*19*18*17*16*15*14*13) = 0.013